OpenStudy (ijlal):

Hey guys ! :) Need help with a new topic called Gamma Functions just started studying these and i am having trouble on to graph them, can someone teach me how to graph Gamma Functions?

OpenStudy (supersmart1001):

www.sosmath.com/calculus/improper/gamma/gamma.html

OpenStudy (ijlal):

yes i know this link @supersmart1001 but it doesn't teach on how to graph them :/

OpenStudy (supersmart1001):

oh then i got to look again...

OpenStudy (ijlal):

\[\Gamma(p)=\int\limits_{0}^{\infty}e^{-x}x^{p-1}dx\]

OpenStudy (ijlal):

for p>0 i need to know how to graph these functions that look like these

OpenStudy (ijlal):

and can some provide some basics on what \[\Gamma(p)=?\] is ? what is Gamma(p) equal to?

OpenStudy (supersmart1001):

i never came to these problems before is this collage???im only in middle

OpenStudy (ijlal):

haha @supersmart1001 no it's higher than college bro :P thanks for trying though :)

OpenStudy (supersmart1001):

ok np

OpenStudy (ijlal):

anyone pls :/ !

OpenStudy (supersmart1001):

yeah

OpenStudy (supersmart1001):

hold on

OpenStudy (supersmart1001):

@SamandDean1967 can u help

OpenStudy (supersmart1001):

@Addicted2Anime can u help

OpenStudy (addicted2anime):

no clue... sorry

OpenStudy (supersmart1001):

ok

OpenStudy (samanddean1967):

let me read up on them see if I understand

OpenStudy (samanddean1967):

@mathstudent55 <<he is good at explaining

OpenStudy (ijlal):

@welshfella @Photon336 guys can you help me out with this one ? or do you know some one who can ? really need to understand this but i am in no rush :)

OpenStudy (photon336):

@hartnn

OpenStudy (ijlal):

@Photon336 thank you :) hopefully he can help me :)

OpenStudy (supersmart1001):

@Summer_Collins @jpharper12345678cars @Atsie @HeatherRunninIt2016 @phi HELP

OpenStudy (supersmart1001):

ok ijlad tell them ur question

OpenStudy (jpharper12345678cars):

ok lets see what going on here

OpenStudy (jpharper12345678cars):

were is the question?

OpenStudy (ijlal):

guys i need to know ho to graph them that's all no further details as i just started learning this topic

OpenStudy (ijlal):

@jpharper12345678cars look above

OpenStudy (ijlal):

@jpharper12345678cars 4th comment from the start graph this function under p>0 conditions

OpenStudy (jpharper12345678cars):

ok

OpenStudy (supersmart1001):

wow all of those people r gone exept jpharper12345678cars

OpenStudy (jpharper12345678cars):

be cause if i can help i dont mind at all

OpenStudy (jpharper12345678cars):

even if i have to do research myself

OpenStudy (jpharper12345678cars):

ok im sorry guy but i need more info on this i need to know what answer your looking for i need help myself

OpenStudy (ijlal):

@jpharper12345678cars its alright :) thanks for trying though

OpenStudy (jpharper12345678cars):

what exactly are you looking for tho

OpenStudy (jpharper12345678cars):

@ShalZayden1234White

OpenStudy (ijlal):

i am looking to graph this function @jpharper12345678cars \[\Gamma(p)=\int\limits_{0}^{\infty}e^{-x} x^{p-1} dx\]

OpenStudy (ijlal):

@Nnesha

OpenStudy (bobo-i-bo):

Do you know properties of the gamma function?

OpenStudy (bobo-i-bo):

ie, for \(p=n, \Gamma(p)=(n-1)!\). So plot those points and then interpolate and then you're done :)